convert the following into degrees, minutes and seconds.
a.) -1.4
b.) 9.6

Answer by MathLover1(20850)


a.) -1.4=-1+(-.4)
This angle is negative.
Degrees = 1
Minutes = (.4)*60=24....we can take away 1 minute and have seconds too
Minutes =23‘......1min=60‘’
Seconds = 60‘’

b.) 9.6
This angle is positive.
Degrees = 9
Minutes =.6*60= 36‘........take away 1 minute and have seconds
Minutes = 35‘
Seconds = 60‘’